Circular No. 8806

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2007 C2 (CATALINA) An apparent asteroidal object discovered by the Catalina Sky Survey

(discovery observation tabulated below), and posted on the Minor Planet Center’s ‘NEOCP’ webpage, has been reported by other observers to show cometary appearance. CCD images taken by H.-C. Lin, C.-S. Lin, and Q.-z. Ye (Lulin Observatory 0.41-m f/8.8 Ritchey-Chrétien telescope) on Feb. 10.55 UT show a small 3′′ coma with an apparent tail around 5′′–6′′ long in p.a. 70^{o}. G. Sostero and E. Guido (Remanzacco, Italy, 0.45-m f/4.4 reflector) report that a co-addition of ten unfiltered 180-s CCD exposures on Feb. 11.8 reveals the possible presence of a broad and faint tail almost 12′′ toward p.a. 340^{o}. D. T. Durig and W. R. O’Farrell (Sewanee, TN, 0.30-m f/4 Schmidt-Cassegrain telescope) write that their CCD images from Feb. 12.09–12.16 show a possible coma of diameter 1′′.5-3′′, asymmetric to the north, with a possible tail 7′′-9′′ long in p.a. 255^{o}.

2007 UT  α_{2000}  δ_{2000}  Mag.
Feb. 9.13925  3^{h}00^{m}14.^{s}90  +22^{o}06′59′′.9  18.7

The available astrometry (including numerous prediscovery observations, the oldest being made at Catalina back to 2006 Oct. 19), the following elliptical orbital elements, and an ephemeris appear on MPEC 2007-C54.

Epoch = 2007 Sept. 17.0 TT

T = 2007 Sept. 4.5278 TT  ω = 179.^{o}4398
e = 0.462035  Ω = 276.1167 }2000.0
q = 3.780116 AU  i = 8.6748

a = 7.026700 AU n^{o} = 0.0529148 P = 18.626 years

COMET P/2007 B3 (GARRADD) Additional observations show that this comet (cf. IAUC 8801) is of

short period; orbital elements from MPEC 2007-C37:

T = 2008 Apr. 8.305 TT  ω = 281.^{o}119
e = 0.11751  Ω = 327.829 }2000.0
q = 4.24543 AU  i = 28.825
a = 4.81071 AU n^{o} = 0.093409 P = 10.55 years
2007 February 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green